OpenStudy (anonymous):

if ( 2 , 8) is a point in a line whose equation is y=-3x+k , determine the value of k? -2 2 22 26

OpenStudy (anonymous):

replace x by 2 and y by 8 and solve for k

OpenStudy (anonymous):

\[8=-3\times 2+k\] \[8=-6+k\] \[14=k\]

OpenStudy (anonymous):

which doesn't seem to be any of your answers, so let me check my work

OpenStudy (anonymous):

no i like my answer for sure. are you sure these are right?

OpenStudy (anonymous):

yeaa i just now noticed i meant to put a negative 8

OpenStudy (anonymous):

then you have \[-8=-6+k\] \[-2=k\]